### v1

# The wanted extension of Fujii and Tsurumaru’s formula for the spectral radius of the Bell-CHSH operator

Abstract. This paper is motivated by a recent paper of Yuki Fujii and Toyohiro Tsurumaru in which they established a beautiful formula for the spectral radius of the Bell-CHSH operator on finite-dimensional Hilbert spaces. To tackle the operator on infinite-dimensional spaces, they elaborated a method based on appropriate approximation of commutators of infinite-dimensional orthogonal projections by commutators of orthogonal projections on finite-dimensional spaces. We here give a proof of Fujii and Tsurumaru’s original formula that works in all dimensions. We also simplify and correct some arguments employed by Fujii and Tsurumaru.
MSC. Primary: 47A25; Secondary: 20C07, 47B35, 47B47, 47B93
Key words. Two orthogonal projections, Commutator, Spectral radius, Bell-CHSH operator
Let $P_{1},Q_{1}$ and $P_{2},Q_{2}$ be two pairs of orthogonal projections on separable complex Hilbert spaces $V_{1}$ and $V_{2}$, respectively, and put $A_{i}=2P_{i}-I,B_{i}=2Q_{i}-I$. The Bell-CHSH operator is the operator
${\mathcal{B}}=(A_{1}+B_{1})\otimes A_{2}+(A_{1}-B_{1})\otimes B_{2}$
on $V_{1}\otimes V_{2}$. See, e.g., [4, 5, 8, 9]. It is known that the spectral radius of $\mathcal{B}$} can be computed by the formula
$\varrho({\mathcal{B}})=\sqrt{4+\varrho([A_{1},B_{1}])\varrho([A_{2},B_{2}])},$ (1)
where [A,B]=ABBA[A,B]=AB-BA. Since $A_{i},B_{i}]=4[P_{i},Q_{i}]$, we are led to finding an expression for the spectral radius of the commutator of two orthogonal projections.
Thus, let $P,Q$ be two orthogonal projections on a separable complex Hilbert space $V$ and put A=2PI,B=2QIA=2P-I,B=2Q-I.
Theorem 1
We always have $\varrho([P,Q])\leq 1/2$ and $\varrho([A,B])\leq 2$.
Proof. By Halmos’ theorem [7] as quoted in Theorem 1.2 of [2], the commutator $[P,Q]$ is unitarily similar to $(0,0,0,0)\oplus\begin{pmatrix}0&W\\ -W&0\end{pmatrix}$ where $W=\sqrt{H(I-H)}$ with a selfadjoint operator $H$ such that $0\leq H\leq 1$. Taking the determinant of $\begin{pmatrix}-\lambda I&W\\ -W&-\lambda I\end{pmatrix}$, is easily seen that the spectral radius of $[P,Q]$ is equal to the spectral radius of $W$, and the latter is $\sqrt{x(1-x)}$ where $x$ is the point in the spectrum of $H$ that is closest to $1/2$. Consequently, we always have $\varrho([P,Q])\leq 1/2$. Since $[A,B]=4[P,Q]$, we get $\varrho([A,B])\leq 2$.    $\square$
Theorem 2
(Tsirelson’s inequality [10]). We always have $\varrho({\mathcal{B}})\leq 2\sqrt{2}$.
Proof. From Theorem 1 and (1) we infer that $\varrho({\mathcal{B}})\leq\sqrt{4+2\cdot 2}=2\sqrt{2}$.   $\square$
Theorem 3
We have $\varrho([A,B])=\sqrt{\lambda^{2}(4-\lambda^{2})}$ where $\lambda$ in $\sigma(A+B)$, the spectrum of $A+B$, is such that $\lambda^{2}/4$ is closest to $1/2$.
Proof. Let us first assume that $P$ and $Q$ are in general position, which means that the terms $(1,1,0,0)$ and $(1,0,1,0)$ in Halmos’ Theorem 1.2 of [2] are absent. Then $A+B-\lambda I$ is unitarily similar to $\begin{pmatrix}2(I-H)-\lambda I&2W\\ 2W&2(H-I)-\lambda I\end{pmatrix}$. By computing the determinant of this matrix we see that $\lambda$ is in the spectrum of $A+B$ if and only if $1-\lambda^{2}/4$ is in the spectrum of $H$. The spectrum of $W^{2}=H(I-H)$ is therefore the set of
$\left(1-\frac{\lambda^{2}}{4}\right)\left(1-\left(1-\frac{\lambda^{2}}{4}\right)\right)=\frac{\lambda^{2}}{4}\left(1-\frac{\lambda^{2}}{4}\right)$
where $\lambda$ ranges over the spectrum of $A+B$. The spectrum of $[P,Q]$ is the set of $\nu$ for which $\det\begin{pmatrix}-\nu I&W\\ -W&-\nu I\end{pmatrix}=\nu^{2}I+W^{2}$ is not invertible. These values $\nu$ are purely imaginary and we may write $\nu=i\mu$ with real $\mu$. Consequently, the spectrum of $[P,Q]$ is the set of $\nu$ such that $\nu^{2}=\mu^{2}=\frac{\lambda^{2}}{4}\left(1-\frac{\lambda^{2}}{4}\right)$ with $\lambda$ in the spectrum of $A+B$. The maximum of $|\mu|$ is attained at a $\lambda$ for which $\lambda^{2}/4$ is closest to $1/2$. For this $\lambda$, we have
$\varrho([A,B])=4\varrho([P,Q])=4|\mu|=4\sqrt{\frac{\lambda^{2}}{4}\left(1-\frac{\lambda^{2}}{4}\right)}=\sqrt{\lambda^{2}(4-\lambda^{2})},$
as asserted.
If $P,Q$ are not in general position, then $[P,Q]$ is as in the proof of Theorem 1 and hence the spectrum of $[P,Q]$ is the union of $\{0\}$ and the above set of $\mu$ for which $\mu^{2}=\frac{\lambda^{2}}{4}\left(1-\frac{\lambda^{2}}{4}\right)$ with $\lambda$ in the spectrum of $\begin{pmatrix}2(I-H)&2W\\ 2W&2(H-I)\end{pmatrix}$, and we therefore get the same expression for $\varrho([A,B])$.    $\square$
For finite-dimensional spaces $V$, this theorem was established by Fujii and Tsurumaru in [6]. They wrote “when $V$ is infinite-dimensional, it is not known whether or how the equality $\varrho([A,B])=\sqrt{\lambda^{2}(4-\lambda^{2})}$ can be extended, either in the form of an equality or an inequality.” Theorem 3 is the extension they were asking about.
The meta theorem of our paper [3] is that Halmos’ two projections theorem is something like Robert Sheckley’s Answerer: no question about two orthogonal projections will go unanswered, provided the question is not foolish, and questions that are not foolish are those where one already knows most of the answer. Fujii and Tsurumaru also employed Halmos’ theorem, but they did this only for the irreducible $2\times 2$ blocks of the orthogonal decomposition of $[A,B]$ resulting from the unitary representation of the infinite dihedral group ${\mathbf{Z}}_{2}\ast{\mathbf{Z}}_{2}$ on $V$. We here apply Halmos’ theorem in its full generality directly to $[A,B]$, without using an orthogonal decomposition as an intermediate step. Since we knew the result of [6] in advance and thus most of the answer, the present study is a nice illustration of the message we tried to convey in [3].
In [6] it is claimed that the $\lambda$ in the formula $\varrho([A,B])=\sqrt{\lambda^{2}(4-\lambda^{2})}$ is the $\lambda$ closest to $\sqrt{2}$. However, there is a subtlety in the matter: a $\lambda$ closest to $\sqrt{2}$ need not be a $\lambda$ for which $\lambda^{2}/4$ is closest to $1/2$. Indeed, if, for example, $A$ and $B$ are given on ${\mathbf{C}}^{3}$ by the involutive matrices
$A=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-1\end{pmatrix},\quad B=\begin{pmatrix}1&0&0\\ 0&2x-1&\sqrt{x(1-x)}\\ 0&\sqrt{x(1-x)}&1-2x\end{pmatrix}$
with $x=0.01$, then the spectrum $\sigma(A+B)$ equals $\{-2\sqrt{x},2\sqrt{x},2\}=\{-0.2,0.2,2\}$, the point $\lambda=2$ is closest to $\sqrt{2}$, but $\lambda=\pm 0.2$ are the points for which $\lambda^{2}/4$ is closest to $1/2$.
Let $\theta\in(0,\pi)$ and put $s=\sin\theta$ and $c=\cos\theta$. Note that $c^{2}+s^{2}=1$ and thus $\frac{1-c}{s}=\frac{s}{1+c}$ and $\frac{1+c}{s}=\frac{s}{1-c}$. Furthermore, if $\theta<\pi/2$ then $c+s>1$, and if $\theta>\pi/2$ then 1<c+s<1-1<c+s<1. Fujii and Tsurumaru consider a pair of involutions in so-called constant-angle one-shifted form. Such a pair is given by
[…]
In [6], it is claimed that $\sigma(A(\theta)+B(\theta))=[-2s,2s]$ for all $\theta\in(0,\pi)$. In fact, the following is true.
Theorem 4
If $\theta\leq\pi/2$ then $\sigma(A(\theta)+B(\theta))=[-2s,2s]\cup\{2\}$, and if $\theta\geq\pi/2$ then $\sigma(A(\theta)+B(\theta))=[-2s,2s]$.
We will give two proofs. They are both based on the observation that $A(\theta)+B(\theta)=T(a)+K$ where $T(a)$ is the infinite tridiagonal Toeplitz matrix with the symbol $a(e^{i\varphi})=s(e^{-i\varphi}+e^{i\varphi})=2s\cos\varphi$ and $K$ is the matrix whose $1,1$ entry is $1+c$ and the other entries of which are zero. We here encounter an instance of a situation in which work with infinite Toeplitz matrices is simpler than working with finite Toeplitz matrices. We note that the appearance of the outlier 22 should not come as too much a surprise. If $\theta=0$, then
$A(0)={\rm diag}(1,-1,1,-1,\ldots),\quad B(0)={\rm diag}(1,1,-1,1,-1,\ldots),$
whence $A(0)+B(0)={\rm diag}(2,0,0,0,0,\ldots)$ with spectrum $\{0,2\}$. On the other hand, if $\theta=\pi$, we get
$A(\pi)={\rm diag}(-1,1,-1,1,\ldots),\quad B(\pi)={\rm diag}(1,-1,1,-1,1,\ldots)$
and thus $A(\pi)+B(\pi)={\rm diag}(0,0,0,0,0,\ldots)$ with spectrum $\{0\}$.
First proof. The essential spectrum of $T(a)+K$ is [2s,2s][-2s,2s], and hence the spectrum of $T(a)+K$ is a subset of $\mathbf{R}$} containing [2s,2s][-2s,2s]. Take $\lambda\in{\mathbf{R}}\setminus[-2s,2s]$. Since $T(a)+K-\lambda I$ is Fredholm of index zero, $\lambda$ is in the spectrum of $T(a)+K$ if and only if the kernel of $T(a)+K-\lambda I$ is not trivial.
Put $\mu=\lambda/s$. The equation $(T(a)+K-\lambda I)x=(T(a)+K-s\mu I)x=0$ is the infinite system
$(1+c-s\mu)x_{0}+sx_{1}=0$
$x_{0}-\mu x_{1}+x_{2}=0$
$x_{1}-\mu x_{2}+x_{3}=0$
$\ldots.$
Suppose this system has a non-trivial solution. Ignoring the first equation, we get $x_{n}=\alpha q_{1}^{n}+\beta q_{2}^{n}$ where $q_{1},q_{2}$ satisfy the equation $q^{2}-\mu q+1=0$. As $q_{1}q_{2}=1$ and $|q_{1}+q_{2}|=|\mu|>2$, we have $|q_{1}|<1$ and $|q_{2}|>1$. Thus, we get $x_{n}=\alpha q_{1}^{n}$. Since $q_{1}-\mu=-1/q_{1}$, the first equation of the infinite system becomes
$0=(1+c-s\mu)\alpha+s\alpha q_{1}=\alpha[1+c+s(q_{1}-\mu)]=\alpha[1+c-s/q_{1}],$
and since $\alpha\neq 0$, it follows that $q_{1}=s/(1+c)$.
We have $s/(1+c)=(1-c)/s\geq 1$ for $\theta\geq\pi/2$, contradicting the requirement that $|q_{1}|<1$. Consequently, for $\theta\geq\pi/2$ no point outside [2s,2s][-2s,2s] is in $\sigma(A(\theta)+B(\theta))$.
So let $\theta<\pi/2$. In that case $q_{1}=s/(1+c)=(1-c)/s$ lies in $(0,1)$, we get $q_{2}=1/q_{1}=(1+c)/s$, and hence
$\frac{\lambda}{s}=\mu=q_{1}+q_{2}=\frac{1+c}{s}+\frac{s}{1+c}=\frac{2(1+c)}{s(1+c)}=\frac{2}{s}.$
It results that $\lambda=2$. For $\lambda=2$ and $x_{n}=\alpha q_{1}^{n}$, the first equation of the system reads
$(c-1)\alpha+s\alpha q_{1}=\alpha[c-1+s(1-c)/s]=0,$
the remaining equations amount to
$1-\frac{2}{s}q_{1}+q_{1}^{2}=1-\frac{2}{s}\frac{s}{1+c}+\frac{1-c}{s}\frac{s}{1+c}=0,$
and as this is true, $\lambda=2$ belongs indeed to the spectrum of $T(a)+K$.    $\square$
Second proof. We know that [2s,2s][-2s,2s] is contained in the spectrum of $T(a)+K$. Take again $\lambda\in{\mathbf{R}}\setminus[-2s,2s]$. The spectrum of $T(a)$ is [2s,2s][-2s,2s] and hence $T(a)-\lambda I=T(a-\lambda)$ is invertible. So we may write
$T(a)+K-\lambda I=T(a-\lambda)[I+T^{-1}(a-\lambda)K],$
and it follows that $T(a)+K-\lambda I$ is invertible if and only if so is $I+T^{-1}(a-\lambda)K$. The matrix of the latter operator is the identity matrix plus $1+c$ times the first column of $T^{-1}(a-\lambda)$, and therefore it is invertible if and only if its $1,1$ entry is nonzero.
To get the inverse of $T(a-\lambda)$, we employ Wiener-Hopf factorization. With $t=e^{i\varphi}$, we have
$a(t)-\lambda=s(t^{-1}+t)-\lambda=st^{-1}\left(1-\frac{\lambda}{s}t+t^{2}\right)$
$=st^{-1}(t-q_{1})(t-q_{2})=-sq_{2}\left(1-\frac{q_{1}}{t}\right)\left(1-\frac{t}{q_{2}}\right),$
where $q_{1},q_{2}$ are as in the first proof. It follows that
$T^{-1}(a-\lambda)=-\frac{1}{sq_{2}}T\left((1-t/q_{2})^{-1}\right)T\left((1-q_{1}/t)^{-1}\right)$
$=-\frac{1}{sq_{2}}\left(\begin{array}[]{cccc}1&&&\\ 1/q_{2}&1&&\\ 1/q_{2}^{2}&1/q_{2}&1&\\ \ldots&\ldots&\ldots&\ldots\end{array}\right)\left(\begin{array}[]{cccc}1&q_{1}&q_{1}^{2}&\ldots\\ &1&q_{1}&\ldots\\ &&1&\ldots\\ \ldots&\ldots&\ldots&\ldots\end{array}\right),$
which shows that the $1,1$ entry of $I+T^{-1}(a-\lambda)K$ is $1-\frac{1+c}{sq_{2}}$. This is zero if and only if $q_{2}=(1+c)/s=s/(1-c)$. As $|q_{2}|>1$ and $s/(1-c)\leq 1$ for $\theta\geq\pi/2$, we conclude that $\lambda$ is not in the spectrum of $T(a)+K$ if $\theta\geq\pi/2$. So let $\theta<\pi/2$. Then $q_{2}>1$ and since $1-\frac{\lambda}{s}q_{2}+q_{2}^{2}=0$, we obtain
$\frac{\lambda}{s}=\frac{1}{q_{2}}(1+q_{2}^{2})=\frac{1}{q_{2}}+q_{2}=\frac{1+c}{s}+\frac{s}{1+c}=\frac{2(1+c)}{s(1+c)}=\frac{2}{s}.$
Thus, necessarily $\lambda=2$. Conversely, if $\lambda=2$, then $q_{2}=(1+c)/s=s/(1-c)$ satisfies the equation $1-\frac{\lambda}{s}q_{2}+q_{2}^{2}=0$ and is greater than $1$. As $1-\frac{1+c}{sq_{2}}=1-\frac{1+c}{s}\frac{s}{1+c}=0$, we see that the $1,1$ entry of $I+T^{-1}(a-\lambda)K$ is zero, and consequently, $\lambda=2$ is in the spectrum of $T(a)+K$.    $\square$
The following theorem is already in [6], but the proof given there is based on the incorrect assumption that $\sigma(A(\theta)+B(\theta))=[-2s,2s]$ for $\theta\in(0,\pi/4)$, which mistake was fortunately neutralized by the fact that the outlier $\lambda=2$, though being closest to $\sqrt{2}$, is not the $\lambda$ for which $\lambda^{2}/4$ is closest to $1/2$.
Theorem 5
In the case of a constant-angle one-shifted form we have
$\varrho([A(\theta),B(\theta)])=\left\{\begin{array}[]{lll}2\sin(2\theta)&\mbox{if}&0<\theta\leq\pi/4,\\ 2&\mbox{if}&\pi/4\leq\theta\leq 3\pi/4,\\ 2|\sin(2\theta)|&\mbox{if}&3\pi/4\leq\theta<\pi.\end{array}\right$.
Proof. We use Theorems 3 and 4. If $\theta\leq\pi/4$, then $\lambda^{2}/4=2^{2}/4=1$ has distance $1/2$ to $1/2$ whereas $\lambda^{2}/4=(2s)^{2}/4=s^{2}$ is at the distance $1/2-s^{2}$ to $1/2$. Consequently, $\lambda=2s$ is a point in the spectrum for which $\lambda^{2}/4$ is closest to $1/2$. This implies that
$\varrho([A(\theta),B(\theta)])=\sqrt{4s^{2}(4-4s^{2})}=4s\sqrt{1-s^{2}}=4sc=2\sin(2\theta).$
If $\pi/4\leq\theta\leq 3\pi/4$, then $\sqrt{2}\in[-2s,2s]$ and hence $\lambda=\sqrt{2}$ minimizes the distance between $\lambda^{2}/4$ and $1/2$. Thus, $\varrho([A(\theta),B(\theta)])=\sqrt{2(4-2)}=2$. Finally, if $3\pi/4\leq\theta<\pi$, then again $\lambda^{2}/4=(2s)^{2}/4=s^{2}$ is closest to $1/2$, and since $c<0$ in this case, we obtain $\varrho([A(\theta),B(\theta)])=2|\sin(2\theta)|$.    $\square$
The general case of two involutions in one-shifted form is given by
$A={\rm diag}\left(r(\omega_{1}),r(\omega_{2}),r(\omega_{3}),\ldots\right),\quad B={\rm diag}\left(1,r(\theta_{1}),r(\theta_{2}),r(\theta_{3}),\ldots\right).$
where $r(\eta)=\begin{pmatrix}\cos\eta&\sin\eta\\ \sin\eta&-\cos\eta\end{pmatrix}$ and $\{\omega_{j}\}_{=1}^{\infty},\{\theta_{j}\}_{=1}^{\infty}$ are sequences of points in $(0,\pi)$. We say that two such involutions are in eventually constant-angle one-shifted form if there is a $j_{0}$ such that $\omega_{j}=\theta_{j}=\theta$ for all $j\geq j_{0}$. Let $P_{n}$ be the projection that sends a sequence $\{x_{j}\}_{j=0}^{\infty}$ to the sequence $\{x_{0},\ldots,x_{n-1},0,0,\ldots\}$. If $T$ is an operator on $\ell^{2}({\mathbf{Z}}_{+})$, then $P_{n}TP_{n}$ may be identified with an $n\times n$ matrix in the natural fashion. The spectrum $\sigma(P_{n}TP_{n})$ is nothing but the set of the eigenvalues of the matrix $P_{n}TP_{n}$.
Theorem 6
If $A$ and $B$ are in eventually constant-angle one-shifted form, then the spectrum of $P_{n}(A+B)P_{n}$ converges in the Hausdorff metric to the spectrum of $A+B$.
Proof. In the case at hand, we have $A+B=T(a)+K$ where, as above, $T(a)$ is the infinite tridiagonal Toeplitz matrix whose symbol is $a(e^{i\varphi})=2s\cos\varphi$ with $s=\sin\theta$ and $K$ is a finite-rank operator with a finite matrix in the upper left corner and zeros elsewhere. By Proposition 2.33 and Theorem  4.16 of [1], the spectrum of $P_{n}(T(a)+K)P_{n}$ converges in the Hausdorff metric to $\sigma(T(a)+K)\cup\sigma(T(a))$. (Note that $T(a)$ is symmetric, which implies that the T(a~)T($\widetilde${a}) occurring in Theorem 4.16 coincides with $T(a)$.) Since $\sigma(T(a))=[-2s,2s]$ and $\sigma(T(a)+K)$ contains the essential spectrum $\sigma_{\rm ess}(T(a))=[-2s,2s]$, we arrive at the assertion.    $\square$
Examples. Let $\omega,\delta\in(0,\pi)$ and
$A={\rm diag}\left(r(\omega),r(\theta),r(\theta),\ldots\right),\quad B={\rm diag}\left(1,r(\theta),r(\theta),r(\theta),\ldots\right).$ (3)
Put $c=\cos\theta,s=\sin\theta,\gamma=\cos\omega,\delta=\sin\omega$. The matrix of $A+B=T(a)+K$ is the tridiagonal Toeplitz matrix generated by $a(e^{i\varphi})=2s\cos\varphi$ with the upper left $2\times 2$ corner matrix $\begin{pmatrix}0&s\\ s&0\end{pmatrix}$ replaced by $\begin{pmatrix}1+\gamma&\delta\\ \delta&c-\gamma\end{pmatrix}$. We know that $2s,2s]\subset\sigma(A+B)$ and that $A+B-\lambda I$ is Fredholm of index zero for $\lambda\notin[-2s,2s]$. Hence, a point $\lambda$ outside [2s,2s][-2s,2s] is in $\sigma(A+B)$ if and only if $A+B$ has a nontrivial kernel in $\ell^{2}({\mathbf{Z}}_{+})$. Since $\|A\|=\|B\|=1$, we also have always $\sigma(A+B)\subset[-2,2]$. Furthermore, if $\pi/4\leq\theta\leq 3\pi/4$, then $2s\geq\sqrt{2}$ and hence $\sqrt{2}\in\sigma(A+B)$, which implies that $\varrho([A,B])=\sqrt{2(4-2)}=2$ in this case.
The equation $(A+B-\lambda I)x=0$ is the system
$(1+\gamma-\lambda)x_{0}+\delta x_{1}=0,$
$\delta x_{0}+(c-\gamma-\lambda)x_{1}+sx_{2}=0,$
$sx_{n}-\lambda x_{n+1}+sx_{n+2}=0\;\>(n\geq 1).$
This is satisfied if and only if $x_{n}=\alpha q^{n-1}$ for $n\geq 1$, where $\alpha\neq 0$ and $q$ is the solution of the equation $q^{2}-(\lambda/s)q+1=0$ whose modulus is less than $1$, and if $x_{0}$ can be chosen so that the first two equations of the system hold. Since $x_{1}=1$ and $x_{2}=q$, the latter is possible if and only if $\delta/(1+\gamma-\lambda)=(c-\gamma-\lambda+sq)/\delta$, that is,
$\lambda^{2}-(c+sq+1)\lambda+(1+\gamma)(c+sq-1)=0.$ (4)
Figure 1: The spectra of $P_{n}(A+B)P_{n}$ multiplied by $i$ for $n=10$ (left) and $n=600$ (right) with $A,B$ as in (3) and with $\omega=\pi/2$ and $\theta=0.1:0.1:3.1$.
Instead of solving this equation, let us use Theorem 6. Figure 1 shows the result for $\omega=\pi/2$ and $\theta=0.1:0.1:3.1$. To be sure that the outliers are not produced by numerical instabilities, let us check whether (4) is true for the rational numbers c=0.8c=-0.8 and $s=0.6$, which corresponds to $\theta=\arccos(-0.8)\approx 2.4981$ and where exact computations can be done by hand. The picture suggests that in this case the two outliers are $\lambda=\pm 1.5$. The roots of $q^{2}-(1.5/0.6)q+1$ are $2$ and $1/2$. Hence, we have to take $q=0.5$. We have $\gamma=0$ and $\delta=1$ for $\omega=\pi/2$. Inserting these values in the left-hand side of (4), we obtain indeed 0. The same happens for $\lambda=-1.5$. Thus, we may conclude that $\sigma(A+B)=[-1.2,1.2]\cup\{-1.5,1.5\}$ for $\omega=\pi/2$ and $\theta\approx 2.4981$. As $1.5^{2}/4-0.5=0.0625<0.14=0.5-0.6^{2}$, we see that in the case at hand the spectral radius $\varrho([A,B])$ is not determined by $s$ but rather by the outliers $\lambda=\pm 1.5$. Theorem 3 gives $\varrho([A,B])=\sqrt{1.5^{2}(4-1.5^{2})}\approx 1.9843$. The value of $\sqrt{4s^{2}(4-4s^{2})}$ is $1.9200$.
Figure 2: Plots of $\lambda_{\max}(\theta)$ where $\lambda_{\max}(\theta)$ is the largest eigenvalue of $P_{n}(A+B)P_{n}$ for $n=100$ with $A,B$ as in (3). In the left picture we took $\omega=\pi/2$, in the right picture we have $\omega=0.1:0.1:3.1$. The red curve is $2\sin\theta$, the green curve is $2|\cos\theta|$. The horizontal lines are at the height $\sqrt{2}$. The rightmost vertical line in the left picture has the abscissa $x_{3}=2.4352$.
The blue curve in the left picture of Figure 2 shows the outliers $\lambda>0$ for $\omega=\pi/2$ and $0<\theta<\pi$. The curve is of constant height $2$ for $0<\theta\leq\pi/2$ and then goes down for $\pi/2<\theta<\pi$. The abscissas of the intersections of the red curve $\theta\mapsto 2\sin\theta$ and the green curve $\theta\mapsto|2\cos\theta|$ are $x_{1}=\pi/4$ and $x_{2}=3\pi/4\approx 2.3562$, and the abscissa of the intersection of the green curve and the blue outlier curve $\theta\mapsto\lambda_{\max}(\theta)$ is $x_{3}\approx 2.4352$. For $\theta>x_{2}$, we have $\lambda>\sqrt{2}$ and $2s<\sqrt{2}$, i.e., both $\lambda^{2}/4-1/2$ and $1/2-s^{2}$ are positive. It follows that $\lambda^{2}/4-1/2<1/2-s^{2}$ if and only if $\lambda^{2}<4(1-s^{2})=4c^{2}$. The latter is the case exactly if the blue curve is below the green curve, that is, for $\theta>x_{3}$. In summary, taking for each $\theta$ the value of $\lambda\in\sigma(A+B)$ for which $\lambda^{2}/4$ is closest to $1/2$, we arrive at the following:
$\varrho([A,B])=\left\{\begin{array}[]{lll}\sqrt{4s^{2}(4-4s^{2})}=2\sin(2\theta)&\mbox{for}&0<\theta\leq x_{1},\\ \sqrt{2(4-2)}=2&\mbox{for}&x_{1}\leq\theta\leq x_{2},\\ \sqrt{4s^{2}(4-4s^{2})}=2|\sin(2\theta)|&\mbox{for}&x_{2}\leq\theta\leq x_{3},\\ \sqrt{\lambda_{\max}(\theta)^{2}(4-\lambda_{\max}(\theta)^{2})}&\mbox{for}&x_{3}\leq\theta<\pi.\end{array}\right$.
The right picture of Figure 2 shows the situation for $\omega=0.1:0.1:3.1$. Figure 3 is for
$A={\rm diag}(r(3/2),r(2),r(\theta),r(\theta),\ldots),\quad B={\rm diag}(1,r(5/2),r(\theta),r(\theta),\ldots).$ (5)
Figure 3: The spectra of $P_{n}(A+B)P_{n}$ multiplied by $i$ for $n=100$ with $A,B$ as in (5) and $\theta=0.1:0.1:3.1$.