### v3

# The wanted extension of Fujii and Tsurumaru’s formula for the spectral radius of the Bell-CHSH operator

Abstract.
This paper is motivated by a recent paper of Yuki Fujii and Toyohiro Tsurumaru in which they established a beautiful formula for the spectral radius of the Bell-CHSH operator on finite-dimensional Hilbert spaces. To tackle the operator on infinite-dimensional spaces, they elaborated a method based on appropriate approximation of commutators of infinite-dimensional orthogonal projections by commutators of orthogonal projections on finite-dimensional spaces. We here give a proof of Fujii and Tsurumaru’s original formula that works in all dimensions. We also present an alternative approximation procedure, uncover the connection of the problem with block Toeplitz operators, and derive good estimates and explicit expressions for the spectral radius in concrete cases.
Key words and phrases:
Two orthogonal projections, Commutator, Spectral radius, Bell-CHSH operator
1991 Mathematics Subject Classification:
Primary 47A25; Secondary 20C07, 47B35, 47B47, 47B93
1. The Bell-CHSH operator
Let $P_{1},Q_{1}$ and $P_{2},Q_{2}$ be two pairs of orthogonal projections on separable complex Hilbert spaces $V_{1}$ and $V_{2}$, respectively, and put $A_{i}=2P_{i}-I,B_{i}=2Q_{i}-I$. The Bell-CHSH operator is the operator
${\mathcal{B}}=(A_{1}+B_{1})\otimes A_{2}+(A_{1}-B_{1})\otimes B_{2}$
on $V_{1}\otimes V_{2}$. The spectral radius $\varrho({\mathcal{B}})$ is of interest in quantum mecanics. See, e.g., [4, 5, 6, 7, 9, 10, 12]. Taking into account that $A_{i}$ and $B_{i}$ are involutions, $A_{i}^{2}=B_{i}^{2}=I$, a simple straightforward computation gives
${\mathcal{B}}^{2}=4(I\otimes I)+[A_{1},B_{1}]\otimes[B_{2},A_{2}],$
where [A,B]=ABBA[A,B]=AB-BA. Consequently, since $\varrho(S)=\varrho(-S)$ for every operator $S$, we get the well-known formula
$\varrho({\mathcal{B}})=\sqrt{4+\varrho([A_{1},B_{1}])\varrho([A_{2},B_{2}])},$ (1)
As $A_{i},B_{i}]=4[P_{i},Q_{i}]$, we are so led to the study of the spectral radius of the commutator of two selfadjoint involutions or equivalently, of two orthogonal projections.
Thus, let $P,Q$ be two orthogonal projections on a separable complex Hilbert space $V$ and put A=2PI,B=2QIA=2P-I,B=2Q-I. We have $\|A\|^{2}=\|AA^{*}\|=\|A^{2}\|=1$ and, analogously, $\|B\|=1$. It follows that
$\varrho([A,B])\leq\|[A,B]\|\leq\|A\|\,\|B\|+\|B\|\,\|A\|=2,$
and hence (1) implies that $\varrho({\mathcal{B}})\leq\sqrt{4+2\cdot 2}=2\sqrt{2}$, which is Tsirelson’s famous inequality [12]. We here are interested in expressions for $\varrho([P,Q])$ and thus for $\varrho([A,B])=4\varrho([P,Q])$ and $\varrho({\mathcal{B}})$ that allow us the exact computation of these quantities for concrete operators $P,Q$.
2. An exact formula for the spectral radius
For a bounded linear operator $T$, we denote by $\sigma(T)$ the spectrum of $T$.
Theorem 1.
We have
$\varrho([A,B])=\max_{\lambda\in\sigma(A+B)}\sqrt{\lambda^{2}(4-\lambda^{2})}=\sqrt{\lambda_{0}^{2}(4-\lambda_{0}^{2})},$
where $\lambda_{0}$ in $\sigma(A+B)$ is a point for which $\lambda^{2}$ is closest to $2$.
Proof. We use Halmos’ theorem [8] as quoted in Theorem 1.2 of [2]. Let us first assume that $P$ and $Q$ are in general position, which means that the terms $(1,1,0,0)$ and $(1,0,1,0)$ in Theorem 1.2 of [2] are absent. Then $A+B-\lambda I$ is unitarily similar to $\begin{pmatrix}2(I-H)-\lambda I&2W\\ 2W&2(H-I)-\lambda I\end{pmatrix}$. By computing the determinant of this matrix we see that $\lambda$ is in the spectrum of $A+B$ if and only if $1-\lambda^{2}/4$ is in the spectrum of $H$. The spectrum of $W^{2}=H(I-H)$ is therefore the set of
$\left(1-\frac{\lambda^{2}}{4}\right)\left(1-\left(1-\frac{\lambda^{2}}{4}\right)\right)=\frac{\lambda^{2}}{4}\left(1-\frac{\lambda^{2}}{4}\right)$
where $\lambda$ ranges over the spectrum of $A+B$. The spectrum of $[P,Q]$ is the set of $\nu$ for which $\det\begin{pmatrix}-\nu I&W\\ -W&-\nu I\end{pmatrix}=\nu^{2}I+W^{2}$ is not invertible. These values $\nu$ are purely imaginary and we may write $\nu=i\mu$ with real $\mu$. Consequently, the spectrum of $[P,Q]$ is the set of $\nu$ such that $\nu^{2}=\mu^{2}=\frac{\lambda^{2}}{4}\left(1-\frac{\lambda^{2}}{4}\right)$ with $\lambda$ in the spectrum of $A+B$. The maximum of $|\mu|$ is attained at a $\lambda_{0}$ for which $\lambda^{2}/4$ is closest to $1/2$ or equivalently, at a $\lambda_{0}$ for which $\lambda^{2}$ is closest to $2$. For this $\lambda_{0}$, we have
$\varrho([A,B])=4\varrho([P,Q])=4|\mu|=4\sqrt{\frac{\lambda_{0}^{2}}{4}\left(1-\frac{\lambda_{0}^{2}}{4}\right)}=\sqrt{\lambda_{0}^{2}(4-\lambda_{0}^{2})},$
as asserted.
If $P,Q$ are not in general position, then $[P,Q]$ is unitarily similar to
$(0,0,0,0)\oplus\begin{pmatrix}0&W\\ -W&0\end{pmatrix}$ (2)
and hence the spectrum of $[P,Q]$ is the union of $\{0\}$ and the above set of $\nu$ for which $\nu^{2}=\frac{\lambda^{2}}{4}\left(1-\frac{\lambda^{2}}{4}\right)$ with $\lambda$ in the spectrum of $\begin{pmatrix}2(I-H)&2W\\ 2W&2(H-I)\end{pmatrix}$, and we therefore get the same expression for $\varrho([A,B])$.    $\square$
For finite-dimensional spaces $V$, this theorem was established by Fujii and Tsurumaru in [6]. They wrote “when $V$ is infinite-dimensional, it is not known whether or how the equality $\varrho([A,B])=\sqrt{\lambda_{0}^{2}(4-\lambda_{0}^{2})}$ can be extended, either in the form of an equality or an inequality.” Theorem 1 is the extension they were asking about.
The meta theorem of our paper [3] is that Halmos’ two projections theorem is something like Robert Sheckley’s Answerer: no question about two orthogonal projections will go unanswered, provided the question is not foolish, and questions that are not foolish are those where one already knows most of the answer. Fujii and Tsurumaru also employed Halmos’ theorem, but they did this only for the irreducible $2\times 2$ blocks of the orthogonal decomposition of $[A,B]$ resulting from the unitary representation of the infinite dihedral group ${\mathbf{Z}}_{2}\ast{\mathbf{Z}}_{2}$ on $V$. We here apply Halmos’ theorem in its full generality directly to $[A,B]$, without using an orthogonal decomposition as an intermediate step. Since we knew the result of [6] in advance and thus most of the answer, the present study is a nice illustration of the message we tried to convey in [3].
Incidentally, and this is again in the vein of [3], one may also employ Halmos’ theorem in its full generality to prove Tsirelson’s inequality. Indeed, we know that $[P,Q]$ is unitarily similar to the operator (2), it is easily seen that the spectral radius of this operator is equal to the spectral radius of $W$, and the latter is $\sqrt{x(1-x)}$ where $x$ is the point in the spectrum of $H$ that is closest to $1/2$. Consequently, we always have $\varrho([P,Q])\leq 1/2$, whence $\varrho([A,B])\leq 2$ and $\varrho({\mathcal{B}})\leq 2\sqrt{2}$.
In [6] it is claimed that the $\lambda_{0}$ in the formula $\varrho([A,B])=\sqrt{\lambda_{0}^{2}(4-\lambda_{0}^{2})}$ is the $\lambda_{0}$ closest to $\sqrt{2}$. However, there is a subtlety in the matter: a $\lambda_{0}$ closest to $\sqrt{2}$ need not be a $\lambda_{0}$ for which $\lambda^{2}$ is closest to $2$. Indeed, if, for example, $A$ and $B$ are given on ${\mathbf{C}}^{3}$ by the selfadjoint and involutive matrices
$A=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-1\end{pmatrix},\quad B=\begin{pmatrix}1&0&0\\ 0&2x-1&\sqrt{x(1-x)}\\ 0&\sqrt{x(1-x)}&1-2x\end{pmatrix}$
with $x=0.01$, then the spectrum $\sigma(A+B)$ equals $\{-2\sqrt{x},2\sqrt{x},2\}=\{-0.2,0.2,2\}$, the point $\lambda=2$ is closest to $\sqrt{2}$, but $\lambda=\pm 0.2$ are the points for which $\lambda^{2}$ is closest to $2$.
3. Pairs in convergent-angle one-shifted form
Let $\theta\in(0,\pi)$ and put $s=\sin\theta$ and $c=\cos\theta$. Note that $c^{2}+s^{2}=1$ and thus $\frac{1-c}{s}=\frac{s}{1+c}$ and $\frac{1+c}{s}=\frac{s}{1-c}$. Furthermore, if $\theta<\pi/2$ then $c+s>1$, and if $\theta>\pi/2$ then 1<c+s<1-1<c+s<1. Fujii and Tsurumaru consider pairs of involutions in so-called constant-angle one-shifted form. Such a pair is given by
[…]
In [6], it is claimed that $\sigma(A(\theta)+B(\theta))=[-2s,2s]$ for all $\theta\in(0,\pi)$. In fact, the following is true.
Theorem 2.
If $\theta\leq\pi/2$ then $\sigma(A(\theta)+B(\theta))=[-2s,2s]\cup\{2\}$, and if $\theta\geq\pi/2$ then $\sigma(A(\theta)+B(\theta))=[-2s,2s]$.
We will give two proofs. They are both based on the observation that $A(\theta)+B(\theta)=T(a)+K$ where $T(a)$ is the infinite tridiagonal Toeplitz matrix with the symbol $a(e^{i\varphi})=s(e^{-i\varphi}+e^{i\varphi})=2s\cos\varphi$ and $K$ is the matrix whose $1,1$ entry is $1+c$ and the other entries of which are zero. We here encounter an instance of a situation in which work with infinite Toeplitz matrices is simpler than working with finite Toeplitz matrices. We note that the appearance of the outlier 22 should not come as too much a surprise. If $\theta=0$, then
$A(0)={\rm diag}(1,-1,1,-1,\ldots),\quad B(0)={\rm diag}(1,1,-1,1,-1,\ldots),$
whence $A(0)+B(0)={\rm diag}(2,0,0,0,0,\ldots)$ with spectrum $\{0,2\}$. On the other hand, if $\theta=\pi$, we get
$A(\pi)={\rm diag}(-1,1,-1,1,\ldots),\quad B(\pi)={\rm diag}(1,-1,1,-1,1,\ldots)$
and thus $A(\pi)+B(\pi)={\rm diag}(0,0,0,0,0,\ldots)$ with spectrum $\{0\}$.
The essential spectrum $\sigma_{\rm ess}(T)$ is the set of $\lambda\in{\mathbf{C}}$ for which $T-\lambda I$ is not Fredholm, that is, not invertible modulo compact operators.
First proof. The essential spectrum of $T(a)+K$ is [2s,2s][-2s,2s], and hence the spectrum of $T(a)+K$ is a subset of $\mathbf{R}$} containing [2s,2s][-2s,2s]. Take $\lambda\in{\mathbf{R}}\setminus[-2s,2s]$. Since $T(a)+K-\lambda I$ is Fredholm of index zero, $\lambda$ is in the spectrum of $T(a)+K$ if and only if the kernel of $T(a)+K-\lambda I$ is not trivial.
Put $\mu=\lambda/s$. The equation $(T(a)+K-\lambda I)x=(T(a)+K-s\mu I)x=0$ is the infinite system
$(1+c-s\mu)x_{0}+sx_{1}=0$
$x_{0}-\mu x_{1}+x_{2}=0$
$x_{1}-\mu x_{2}+x_{3}=0$
$\ldots.$
Suppose this system has a non-trivial solution. Ignoring the first equation, we get $x_{n}=\alpha q_{1}^{n}+\beta q_{2}^{n}$ where $q_{1},q_{2}$ satisfy the equation $q^{2}-\mu q+1=0$. As $q_{1}q_{2}=1$ and $|q_{1}+q_{2}|=|\mu|>2$, we have $|q_{1}|<1$ and $|q_{2}|>1$. Thus, we get $x_{n}=\alpha q_{1}^{n}$. Since $q_{1}-\mu=-1/q_{1}$, the first equation of the infinite system becomes
$0=(1+c-s\mu)\alpha+s\alpha q_{1}=\alpha[1+c+s(q_{1}-\mu)]=\alpha[1+c-s/q_{1}],$
and since $\alpha\neq 0$, it follows that $q_{1}=s/(1+c)$.
We have $s/(1+c)=(1-c)/s\geq 1$ for $\theta\geq\pi/2$, contradicting the requirement that $|q_{1}|<1$. Consequently, for $\theta\geq\pi/2$ no point outside [2s,2s][-2s,2s] is in $\sigma(A(\theta)+B(\theta))$.
So let $\theta<\pi/2$. In that case $q_{1}=s/(1+c)=(1-c)/s$ lies in $(0,1)$, we get $q_{2}=1/q_{1}=(1+c)/s$, and hence
$\frac{\lambda}{s}=\mu=q_{1}+q_{2}=\frac{1+c}{s}+\frac{s}{1+c}=\frac{2(1+c)}{s(1+c)}=\frac{2}{s}.$
It results that $\lambda=2$. For $\lambda=2$ and $x_{n}=\alpha q_{1}^{n}$, the first equation of the system reads
$(c-1)\alpha+s\alpha q_{1}=\alpha[c-1+s(1-c)/s]=0,$
the remaining equations amount to
$1-\frac{2}{s}q_{1}+q_{1}^{2}=1-\frac{2}{s}\frac{s}{1+c}+\frac{1-c}{s}\frac{s}{1+c}=0,$
and as this is true, $\lambda=2$ belongs indeed to the spectrum of $T(a)+K$.    $\square$
Second proof. We know that [2s,2s][-2s,2s] is contained in the spectrum of $T(a)+K$. Take again $\lambda\in{\mathbf{R}}\setminus[-2s,2s]$. The spectrum of $T(a)$ is [2s,2s][-2s,2s] and hence $T(a)-\lambda I=T(a-\lambda)$ is invertible. So we may write
$T(a)+K-\lambda I=T(a-\lambda)[I+T^{-1}(a-\lambda)K],$
and it follows that $T(a)+K-\lambda I$ is invertible if and only if so is $I+T^{-1}(a-\lambda)K$. The matrix of the latter operator is the identity matrix plus $1+c$ times the first column of $T^{-1}(a-\lambda)$, and therefore it is invertible if and only if its $1,1$ entry is nonzero.
To get the inverse of $T(a-\lambda)$, we employ Wiener-Hopf factorization. With $t=e^{i\varphi}$, we have
$a(t)-\lambda=s(t^{-1}+t)-\lambda=st^{-1}\left(1-\frac{\lambda}{s}t+t^{2}\right)$
$=st^{-1}(t-q_{1})(t-q_{2})=-sq_{2}\left(1-\frac{q_{1}}{t}\right)\left(1-\frac{t}{q_{2}}\right),$
where $q_{1},q_{2}$ are as in the first proof. It follows that
$T^{-1}(a-\lambda)=-\frac{1}{sq_{2}}T\left((1-t/q_{2})^{-1}\right)T\left((1-q_{1}/t)^{-1}\right)$
$=-\frac{1}{sq_{2}}\left(\begin{array}[]{cccc}1&&&\\ 1/q_{2}&1&&\\ 1/q_{2}^{2}&1/q_{2}&1&\\ \ldots&\ldots&\ldots&\ldots\end{array}\right)\left(\begin{array}[]{cccc}1&q_{1}&q_{1}^{2}&\ldots\\ &1&q_{1}&\ldots\\ &&1&\ldots\\ \ldots&\ldots&\ldots&\ldots\end{array}\right),$
which shows that the $1,1$ entry of $I+T^{-1}(a-\lambda)K$ is $1-\frac{1+c}{sq_{2}}$. This is zero if and only if $q_{2}=(1+c)/s=s/(1-c)$. As $|q_{2}|>1$ and $s/(1-c)\leq 1$ for $\theta\geq\pi/2$, we conclude that $\lambda$ is not in the spectrum of $T(a)+K$ if $\theta\geq\pi/2$. So let $\theta<\pi/2$. Then $q_{2}>1$ and since $1-\frac{\lambda}{s}q_{2}+q_{2}^{2}=0$, we obtain
$\frac{\lambda}{s}=\frac{1}{q_{2}}(1+q_{2}^{2})=\frac{1}{q_{2}}+q_{2}=\frac{1+c}{s}+\frac{s}{1+c}=\frac{2(1+c)}{s(1+c)}=\frac{2}{s}.$
Thus, necessarily $\lambda=2$. Conversely, if $\lambda=2$, then $q_{2}=(1+c)/s=s/(1-c)$ satisfies the equation $1-\frac{\lambda}{s}q_{2}+q_{2}^{2}=0$ and is greater than $1$. As $1-\frac{1+c}{sq_{2}}=1-\frac{1+c}{s}\frac{s}{1+c}=0$, we see that the $1,1$ entry of $I+T^{-1}(a-\lambda)K$ is zero, and consequently, $\lambda=2$ is in the spectrum of $T(a)+K$.    $\square$
The following theorem is already in [6], but the proof given there is based on the incorrect assumption that $\sigma(A(\theta)+B(\theta))=[-2s,2s]$ for $\theta\in(0,\pi/4)$, which mistake was fortunately neutralized by the fact that the outlier $\lambda=2$, though being closest to $\sqrt{2}$, is not the $\lambda$ for which $\lambda^{2}$ is closest to $2$.
Theorem 3.
In the case of a constant-angle one-shifted form we have
$\varrho([A(\theta),B(\theta)])=\left\{\begin{array}[]{lll}2\sin(2\theta)&\mbox{if}&0<\theta\leq\pi/4,\\ 2&\mbox{if}&\pi/4\leq\theta\leq 3\pi/4,\\ 2|\sin(2\theta)|&\mbox{if}&3\pi/4\leq\theta<\pi.\end{array}\right$.
Proof. We use Theorems 1 and 2. If $\theta\leq\pi/4$, then $\lambda^{2}=2^{2}=4$ has distance $2$ to $2$ whereas $\lambda^{2}=(2s)^{2}=4s^{2}$ is at the distance $2-4s^{2}$ to $2$. Consequently, $\lambda_{0}=2s$ is a point in the spectrum for which $\lambda^{2}$ is closest to $2$. This implies that
$\varrho([A(\theta),B(\theta)])=\sqrt{4s^{2}(4-4s^{2})}=4s\sqrt{1-s^{2}}=4sc=2\sin(2\theta).$
If $\pi/4\leq\theta\leq 3\pi/4$, then $\sqrt{2}\in[-2s,2s]$ and hence $\lambda_{0}=\sqrt{2}$ minimizes the distance between $\lambda^{2}$ and $2$. Thus, $\varrho([A(\theta),B(\theta)])=\sqrt{2(4-2)}=2$. Finally, if $3\pi/4\leq\theta<\pi$, then again $\lambda_{0}^{2}=(2s)^{2}=4s^{2}$ is closest to $2$, and since $c<0$ in this case, we obtain $\varrho([A(\theta),B(\theta)])=2|\sin(2\theta)|$.    $\square$
The general case of two involutions in one-shifted form is given by
$A={\rm diag}\left(r(\omega_{1}),r(\omega_{2}),r(\omega_{3}),\ldots\right),\quad B={\rm diag}\left(1,r(\theta_{1}),r(\theta_{2}),r(\theta_{3}),\ldots\right).$
where $r(\eta)=\begin{pmatrix}\cos\eta&\sin\eta\\ \sin\eta&-\cos\eta\end{pmatrix}$ and $\{\omega_{j}\}_{=1}^{\infty},\{\theta_{j}\}_{=1}^{\infty}$ are sequences of points in $(0,\pi)$. We say that two such involutions are in convergent-angle one-shifted form if there is a $\theta\in(0,\pi)$ such that $\omega_{j}\to\theta$ and $\theta_{j}\to\theta$ as $j\to\infty$. Let $P_{n}$ be the projection that sends a sequence $\{x_{j}\}_{j=0}^{\infty}$ to the sequence $\{x_{0},\ldots,x_{n-1},0,0,\ldots\}$. If $T$ is an operator on $\ell^{2}({\mathbf{Z}}_{+})$, then $P_{n}TP_{n}$ may be identified with an $n\times n$ matrix in the natural fashion. The spectrum $\sigma(P_{n}TP_{n})$ is nothing but the set of the eigenvalues of the matrix $P_{n}TP_{n}$.
Theorem 4.
If $A$ and $B$ are in convergent-angle one-shifted form, then the spectrum of $P_{n}(A+B)P_{n}$ converges in the Hausdorff metric to the spectrum of $A+B$.
Proof. In the case at hand, we have $A+B=T(a)+K$ where, as above, $T(a)$ is the infinite tridiagonal Toeplitz matrix whose symbol is $a(e^{i\varphi})=2s\cos\varphi$ with $s=\sin\theta$ and $K$ is a compact operator (given by a tridiagonal matrix whose entries go to zero along each diagonal). By Proposition 2.33 and Theorem  4.16 of [1], the spectrum of $P_{n}(T(a)+K)P_{n}$ converges in the Hausdorff metric to $\sigma(T(a)+K)\cup\sigma(T(a))$. (Note that $T(a)$ is symmetric, which implies that the T(a~)T($\widetilde${a}) occurring in Theorem 4.16 coincides with $T(a)$.) Since $\sigma(T(a))=[-2s,2s]$ and $\sigma(T(a)+K)$ contains the essential spectrum $\sigma_{\rm ess}(T(a))=[-2s,2s]$, we arrive at the assertion.    $\square$
Examples. Let $\omega,\delta\in(0,\pi)$ and
$A={\rm diag}\left(r(\omega),r(\theta),r(\theta),\ldots\right),\quad B={\rm diag}\left(1,r(\theta),r(\theta),r(\theta),\ldots\right).$ (4)
Put $c=\cos\theta,s=\sin\theta,\gamma=\cos\omega,\delta=\sin\omega$. The matrix of $A+B=T(a)+K$ is the tridiagonal Toeplitz matrix generated by $a(e^{i\varphi})=2s\cos\varphi$ with the upper left $2\times 2$ corner matrix $\begin{pmatrix}0&s\\ s&0\end{pmatrix}$ replaced by $\begin{pmatrix}1+\gamma&\delta\\ \delta&c-\gamma\end{pmatrix}$. We know that $2s,2s]\subset\sigma(A+B)$ and that $A+B-\lambda I$ is Fredholm of index zero for $\lambda\notin[-2s,2s]$. Hence, a point $\lambda$ outside [2s,2s][-2s,2s] is in $\sigma(A+B)$ if and only if $A+B$ has a nontrivial kernel in $\ell^{2}({\mathbf{Z}}_{+})$. Since $\|A\|=\|B\|=1$, we also have always $\sigma(A+B)\subset[-2,2]$. Furthermore, if $\pi/4\leq\theta\leq 3\pi/4$, then $2s\geq\sqrt{2}$ and hence $\sqrt{2}\in\sigma(A+B)$, which implies that $\varrho([A,B])=\sqrt{2(4-2)}=2$ in this case.
The equation $(A+B-\lambda I)x=0$ is the system
$(1+\gamma-\lambda)x_{0}+\delta x_{1}=0,$
$\delta x_{0}+(c-\gamma-\lambda)x_{1}+sx_{2}=0,$
$sx_{n}-\lambda x_{n+1}+sx_{n+2}=0\;\>(n\geq 1).$
This is satisfied if and only if $x_{n}=\alpha q^{n-1}$ for $n\geq 1$, where $\alpha\neq 0$ and $q$ is the solution of the equation $q^{2}-(\lambda/s)q+1=0$ whose modulus is less than $1$, and if $x_{0}$ can be chosen so that the first two equations of the system hold. Since $x_{1}=1$ and $x_{2}=q$, the latter is possible if and only if $\delta/(1+\gamma-\lambda)=(c-\gamma-\lambda+sq)/\delta$, that is,
$\lambda^{2}-(c+sq+1)\lambda+(1+\gamma)(c+sq-1)=0.$ (5)
Figure 1. The spectra of $P_{n}(A+B)P_{n}$ multiplied by $i$ for $n=10$ (left) and $n=600$ (right) with $A,B$ as in (3) and with $\omega=\pi/2$ and $\theta=0.1:0.1:3.1$. The solid curves are plots of $\pm 2\sin\theta$.
Figure 2. Plots of $\lambda_{\max}(\theta)$ where $\lambda_{\max}(\theta)$ is the largest eigenvalue of $P_{n}(A+B)P_{n}$ for $n=100$ with $A,B$ as in (3). In the left picture we took $\omega=\pi/2$, in the right picture we have $\omega=0.1:0.1:3.1$. The red curve is $2\sin\theta$, the green curve is $2|\cos\theta|$. The horizontal lines are at the height $\sqrt{2}$. The rightmost vertical line in the left picture has the abscissa $x_{3}=2.4352$.
Instead of solving this equation, let us use Theorem 4. Figure 1 shows the result for $\omega=\pi/2$ and $\theta=0.1:0.1:3.1$. To be sure that the outliers are not produced by numerical instabilities, let us check whether (4) is true for the rational numbers c=0.8c=-0.8 and $s=0.6$, which corresponds to $\theta=\arccos(-0.8)\approx 2.4981$ and where exact computations can be done by hand. The picture suggests that in this case the two outliers are $\lambda=\pm 1.5$. The roots of $q^{2}-(1.5/0.6)q+1$ are $2$ and $1/2$. Hence, we have to take $q=0.5$. We have $\gamma=0$ and $\delta=1$ for $\omega=\pi/2$. Inserting these values in the left-hand side of (4), we obtain indeed 0. The same happens for $\lambda=-1.5$. Thus, we may conclude that $\sigma(A+B)=[-1.2,1.2]\cup\{-1.5,1.5\}$ for $\omega=\pi/2$ and $\theta\approx 2.4981$. As $1.5^{2}-2=0.25<0.56=2-4\cdot 0.6^{2}$, we see that in the case at hand the spectral radius $\varrho([A,B])$ is not determined by $s$ but rather by the outliers $\lambda=\pm 1.5$. Theorem 1 gives $\varrho([A,B])=\sqrt{1.5^{2}(4-1.5^{2})}\approx 1.9843$. The value of $\sqrt{4s^{2}(4-4s^{2})}$ is $1.9200$.
Figure 3. The spectra of $P_{n}(A+B)P_{n}$ multiplied by $i$ for $n=100$ with $A,B$ as in (5) and $\theta=0.1:0.1:3.1$.
The blue curve in the left picture of Figure 2 shows the outliers $\lambda>0$ for $\omega=\pi/2$ and $0<\theta<\pi$. The curve is of constant height $2$ for $0<\theta\leq\pi/2$ and then goes down for $\pi/2<\theta<\pi$. The abscissas of the intersections of the red curve $\theta\mapsto 2\sin\theta$ and the green curve $\theta\mapsto|2\cos\theta|$ are $x_{1}=\pi/4$ and $x_{2}=3\pi/4\approx 2.3562$, and the abscissa of the intersection of the green curve and the blue outlier curve $\theta\mapsto\lambda_{\max}(\theta)$ is $x_{3}\approx 2.4352$. For $\theta>x_{2}$, we have $\lambda>\sqrt{2}$ and $2s<\sqrt{2}$, i.e., both $\lambda^{2}-2$ and $2-4s^{2}$ are positive. It follows that $\lambda^{2}-2<2-4s^{2}$ if and only if $\lambda^{2}<4(1-s^{2})=4c^{2}$. The latter is the case exactly if the blue curve is below the green curve, that is, for $\theta>x_{3}$. In summary, taking for each $\theta$ the value of $\lambda_{0}\in\sigma(A+B)$ for which $\lambda^{2}$ is closest to $2$, we arrive at the following:
$\varrho([A,B])=\left\{\begin{array}[]{lll}\sqrt{4s^{2}(4-4s^{2})}=2\sin(2\theta)&\mbox{for}&0<\theta\leq x_{1},\\ \sqrt{2(4-2)}=2&\mbox{for}&x_{1}\leq\theta\leq x_{2},\\ \sqrt{4s^{2}(4-4s^{2})}=2|\sin(2\theta)|&\mbox{for}&x_{2}\leq\theta\leq x_{3},\\ \sqrt{\lambda_{\max}(\theta)^{2}(4-\lambda_{\max}(\theta)^{2})}&\mbox{for}&x_{3}\leq\theta<\pi.\end{array}\right$.
The right picture of Figure 2 shows the situation for $\omega=0.1:0.1:3.1$. Figure 3 is for
$A={\rm diag}(r(3/2),r(2),r(\theta),r(\theta),\ldots),B={\rm diag}(1,r(5/2),r(\theta),r(\theta),\ldots).$ (6)
4. Pairs in one-shifted form with two converging angles
A natural generalization of pairs in convergent-angle one-shifted form is pairs of the form
$A={\rm diag}\left(r(\omega_{1}),r(\omega_{2}),r(\omega_{3}),\ldots\right),\quad B={\rm diag}\left(1,r(\theta_{1}),r(\theta_{2}),r(\theta_{3}),\ldots\right)$
with numbers $\omega,\theta\in(0,\pi)$ such that $\omega_{j}\to\omega$ and $\theta_{j}\to\theta$ as $j\to\infty$. Then $A+B=T(a)+K$ with a selfadjoint block Toeplitz matrix $T(a)$ and a compact operator $K$. The diagonal block of $T(a)$ is $\begin{pmatrix}\gamma-c&\delta\\ \delta&c-\gamma\end{pmatrix}$, the block in the subdiagonal equals $\begin{pmatrix}0&s\\ 0&0\end{pmatrix}$, the block in the superdiagonal is equal to $\begin{pmatrix}0&0\\ s&0\end{pmatrix}$, and the other blocks are zero. Consequently,
$a(e^{i\varphi})=\begin{pmatrix}\gamma-c&\delta+se^{i\varphi}\\ \delta+se^{-i\varphi}&c-\gamma\end{pmatrix}$, (7)
where $\gamma=\cos\omega,\delta=\sin\omega,c=\cos\theta,s=\sin\theta$. Put
$\lambda_{1}=\sqrt{(c-\gamma)^{2}+(s-\delta)^{2}}=2\left|\sin\frac{\theta-\omega}{2}\right|,$
$\lambda_{2}=\sqrt{(c-\gamma)^{2}+(s+\delta)^{2}}=2\left|\sin\frac{\theta+\omega}{2}\right|.$
Note that $s$ and $\delta$ are positive. Unless $s=\delta$, which case was disposed of in the previous section, we have
$2\leq-\lambda_{2}<-\lambda_{1}<c-\gamma<\lambda_{1}<\lambda_{2}\leq 2.$ (8)
In cases in which the spectrum of $A+B=T(a)+K$ is available, Theorem 1 gives the spectral radius. To determine this spectrum, it would be desirable to have a result like Theorem 4. Since the blocks of $T(a)$ are of size $2$, we consider only $2n\times 2n$ truncations. It is known that
$\sigma(P_{2n}(T(a)+K)P_{2n})\to\sigma(T(a)+K)\cup\sigma(T({\tilde{a}}))=:\Lambda$ (9)
in the Hausdorff metric; see page 197 of [2]. Here $\tilde{a}$} is given by
${\tilde{a}}(e^{i\varphi})=a(e^{-i\varphi})=\begin{pmatrix}\gamma-c&\delta+se^{-i\varphi}\\ \delta+se^{i\varphi}&c-\gamma\end{pmatrix}$. (10)
Thus, the convergence of $\sigma(P_{2n}(T(a)+K)P_{2n})$ to some limit set $\Lambda$ is guaranteed. However, the appearance of the spectrum $)\sigma(T({\tilde{a}}$)) in (8) causes some trouble since it forces us to find out which parts of the limit set $\Lambda$ constitute $\sigma(T(a)+K)=\sigma(A+B)$ and which parts result only from the contribution of $)\sigma(T({\tilde{a}}$)).
Lemma 5.
Let $a$ be given by (6). The essential spectra of $T(a)$ and T(a~)T($\tilde${a}}) are
$\sigma_{\rm ess}(T(a))=\sigma_{\rm ess}(T({\tilde{a}}))=[-\lambda_{2},-\lambda_{1}]\cup[\lambda_{1},\lambda_{2}]$
and the spectrum of T(a~)T($\tilde${a}}) is $)\sigma_{\rm ess}(T({\tilde{a}}$)) if $s\leq\delta$ and $\sigma_{\rm ess}(T({\tilde{a}}))\cup\{c-\gamma\}$ if $s>\delta$.
Proof. The essential spectrum of $T(a)$ is the set of $\lambda\in{\mathbf{R}}$ for which the determinant of $a(e^{i\varphi})-\lambda I$ vanishes for some $e^{i\varphi}\in{\mathbf{T}}$; see, e.g., Theorem 6.5 of [2]. We have
$\det(a(e^{i\varphi})-\lambda I)=\lambda^{2}-(c-\gamma)^{2}-\delta^{2}-2\delta s\cos\varphi-s^{2}=0$
if and only if $\lambda^{2}=(c-\gamma)^{2}+\delta^{2}+2\delta s\cos\varphi+s^{2}$, and since $2\cos\varphi$ fills the segment [2,2][-2,2], it follows that the essential spectrum of $T(a)$ is equal to $\lambda_{2},-\lambda_{1}]\cup[\lambda_{1},\lambda_{2}]$. As $\cos\varphi=\cos(-\varphi)$, we get the same result for $)\sigma_{\rm ess}(T({\tilde{a}}$)).
That the spectrum $)\sigma(T({\tilde{a}}$)) is a subset of the segment $\lambda_{2},\lambda_{2}]$ can be deduced from general results on positively definite Hermitian matrix function; see Section 7.3 of [11]. However, we have also to look whether or not the spectrum contains points in the gap $(-\lambda_{1},\lambda_{2})$. We therefore proceed as follows.
Take $)\lambda\notin\sigma_{\rm ess}(T({\tilde{a}}$)). Then the selfadjoint operator $T({\tilde{a}})-\lambda I$ is Fredholm of index zero, and hence $\lambda$ belongs to $)\sigma(T({\tilde{a}}$)) if and only if the null space of this operator is nontrivial. The equation
$(T({\tilde{a}})-\lambda I)\begin{pmatrix}f_{+}\\ g_{+}\end{pmatrix}=0$
with $f_{+},g_{+}$ in the Hardy space $H^{2}({\mathbf{T}})$ is equivalent to the existence of functions $h_{-},k_{-}\in\overline{H^{2}}({\mathbf{T}})$ such that $h_{-}(0)=k_{-}(0)=0$ and
$(\gamma-c-\lambda)f_{+}(z)+(\delta+s\overline{z})g_{+}(z)=h_{-}(z),$
$(\delta+s{z})f_{+}(z)+(c-\gamma-\lambda)g_{+}(z)=k_{-}(z)$
for $z\in{\mathbf{T}}$. A moment’s thought reveals that this can only be true for $k_{-}(z)=0$ and $h_{-}(z)=\alpha\overline{z}$ with some constant $\alpha$. Thus, we arrive at the system
$(\gamma-c-\lambda)zf_{+}(z)+(\delta z+s)g_{+}(z)=\alpha,$
$(\delta+sz)f_{+}(z)+(c-\gamma-\lambda)g_{+}(z)=0.$
Let $\lambda\neq c-\gamma$. Solving the second of these equation for $g_{+}(z)$ and inserting the result in the first equation, we get after some straightforward computation the equation $p_{+}(z)f_{+}(z)/(c-\gamma+\lambda)=\alpha$ with
$p_{+}(z)=s\delta\left[1+\frac{2-2c\gamma-\lambda^{2}}{s\delta}z+z^{2}\right].$
This equation has a nontrivial solution $f_{+}\in H^{2}({\mathbf{T}})$ if and only if $1/p_{+}$ is in $H^{2}({\mathbf{T}})$, which in turn happens if and only if both roots of $p_{+}(z)$ lie outside the closed unit disk. As, by Vieta, the product of the two roots is $1$, this is impossible. Consequently, a point $\lambda\neq c-\gamma$ never belongs to the spectrum of T(a~)T($\tilde${a}}).
So consider $\lambda=c-\gamma$. Then the system is reduced to
$2(\gamma-c-\lambda)zf_{+}(z)+(\delta z+s)g_{+}(z)=\alpha,$
$(\delta+sz)f_{+}(z)=0,$
and this holds with a nontrivial function $g_{+}\in H^{2}({\mathbf{T}})$ if and only if $\delta z+s$ has no zero in the closed unit disk, that is, if and only if $s>\delta$.    $\square$
As said, the point $\lambda=c-\gamma$ causes some trouble. We have at least the following.
Theorem 6.
Let $\lambda_{0}$ be a point in $\Lambda$ for which $\lambda^{2}$ is closest to $2$. If $\lambda_{0}\neq c-\gamma$, then $\varrho([A,B])=\sqrt{\lambda_{0}^{2}(4-\lambda_{0}^{2})}$, and if $\lambda_{0}=c-\gamma$, then
$\sqrt{\lambda_{*}^{2}(4-\lambda_{*}^{2})}\leq\varrho([A,B])\leq\sqrt{\lambda_{0}^{2}(4-\lambda_{0}^{2})}$
where $\lambda_{*}$ is a point in $\Lambda\setminus\{c-\gamma\}$ for which $\lambda^{2}$ is closest to $2$.
Proof. The essential spectrum of $T(a)+K$ is equal to the essential spectrum of $T(a)$ and is a subset of $\sigma(T(a)+K)$. From Lemma 5 we therefore see that
$\sigma(T(a)+K)=[-\lambda_{2},-\lambda_{1}]\cup[\lambda_{1},\lambda_{2}]\cup S,$
with some (possibly empty) set $S$. Hence, again by Lemma 5,
$\Lambda=[-\lambda_{2},-\lambda_{1}]\cup[\lambda_{1},\lambda_{2}]\cup S\cup X,$
where $X=\{c-\gamma\}$ if $c-\gamma\in\sigma(T(a))$ and $X=\emptyset$ if $c-\gamma\notin\sigma(T(a))$. Thus, if $\lambda_{0}\neq c-\gamma$, then $\lambda_{0}\in\sigma(T(a)+K)$ and Theorem 1 gives $\varrho([A,B])=\sqrt{\lambda_{0}^{2}(4-\lambda_{0}^{2})}$, while if $\lambda_{0}=c-\gamma$, it may be that $\lambda_{0}$ is not in $\sigma(T(a)+K)$, so that we must exclude $\lambda_{0}$ from the candidates and have to restrict ourselves to $\Lambda$ minus $\{\lambda_{0}\}$, which gives a lower bound for $\varrho([A,B])$.    $\square$
The following theorem contains two constellations in which $c-\gamma$ is ruled out automatically. We remark that
$\sqrt{2}\in(\lambda_{2},2)\;\>\Longleftrightarrow\;\>\theta+\omega\in(0,\pi/2)\cup(3\pi/2,2\pi),$
$\sqrt{2}\in(0,\lambda_{1})\;\>\Longleftrightarrow\;\>\theta-\omega\in(-\pi/2,\pi/2).$
Theorem 7.
Let $\Lambda$ be the limit of $\sigma(P_{2n}(A+B)P_{2n})$ as $n\to\infty$.
(a) If $\sqrt{2}\in(\lambda_{2},2)$, then $\varrho([A,B])=\sqrt{\lambda_{0}^{2}(4-\lambda_{0}^{2})}$ where $\lambda_{0}\in\Lambda$ is a point for which $\lambda^{2}$ is closest to $2$.
(b) If $\sqrt{2}\in[\lambda_{1},\lambda_{2}]$, then $\varrho([A,B])=2$.
(c) If $\sqrt{2}\in(0,\lambda_{1})$, then $\sqrt{\lambda_{1}^{2}(4-\lambda_{1}^{2})}\leq\varrho([A,B])\leq 2$.
Proof. (a) Lemma 5 and (7) tell us that $)\sigma(T({\tilde{a}}$)) is a subset of the segment $\lambda_{2},\lambda_{2}]$ and that the set $\Lambda$ differs from $\sigma(T(a)+K)$ only by the possible point $c-\gamma$ in $)\sigma(T({\tilde{a}}$)), which lies in $\lambda_{1},\lambda_{1}]$. If $c-\gamma=\pm\lambda_{1}$ (which happens only for $s=\delta$ and thus $\lambda_{1}=0$), then $c-\gamma$ belongs to $\sigma(T(a)+K)$. Thus, let $c-\gamma\in(-\lambda_{1},\lambda_{1})$. Since $\lambda_{2}<\sqrt{2}$, we have
$2-\lambda^{2}>2-\lambda_{1}^{2}\geq 2-\lambda_{2}^{2}>0$
for $\lambda\in(-\lambda_{1},\lambda_{1})$, and hence the point $\lambda=c-\gamma$ cannot be a point in $\Lambda$ for which $\lambda^{2}$ is closest to $2$. It follows that the points $\lambda$ in $\Lambda$ for which $\lambda^{2}$ is closest to 22 belong to $\sigma(T(a)+K)$, and so Theorem 1 gives the assertion.
(b) If $\sqrt{2}\in[\lambda_{1},\lambda_{2}]$, then $\lambda=\sqrt{2}$ belongs to $\sigma(T(a)+K)$, it is a point for which $\lambda^{2}$ is closest to $2$, and from Theorem 1 we get $\varrho([A,B])=2$.
(c) In this case it may be that $\lambda=c-\gamma$ is a point for which $\lambda^{2}$ is closest to 22 but that this point lies in $\Lambda\setminus\sigma(T(a)+K)$. Since $\lambda_{1}$ is definitely a point in $\sigma(T(a)+K)$, Theorem 1 implies at least the lower bound $\sqrt{\lambda_{1}^{2}(4-\lambda_{1}^{2})}$ for $\varrho([A,B])$. Finally, we know that always $\varrho([A,B])\leq 2$.    $\square$
5. Pairs in one-shifted form with two constant angles
Let us finally consider the case were both angles are constant,
$A={\rm diag}\left(r(\omega),r(\omega),r(\omega),\ldots\right),\;\>B={\rm diag}\left(1,r(\theta),r(\theta),r(\theta),\ldots\right)$ (11)
with two numbers $\omega,\theta$ in the interval $(0,\pi)$. As above, let
$c=\cos\theta,\;s=\sin\theta,\;\gamma=\cos\omega,\;\delta=\sin\omega.$
In that case $A+B=T(a)+K$ with $a$ as in (6) and with
$K={\rm diag}(1+c,0,0,\ldots).$
Figure 4 shows the limit sets for $\omega=0.3$ and $\omega=2.4$. These figures suggest that $c-\gamma$ belongs to $\sigma(A+B)$ if and only if $\theta=\omega$ or $\theta=\pi-\omega$, which happens if and only if $c=\gamma$ or $c=-\gamma$. If $\theta=\omega$, then indeed $c-\gamma=0=\lambda_{1}$ and if $\theta=\pi-\omega$, then $c-\gamma=2c=\pm\lambda_{1}$, and $\pm\lambda_{1}$ belong to $\sigma(T(a))+K)$. The case $c\neq\pm\gamma$ is settled by the following lemma.
Lemma 8.
If $A,B$ are given by (10) and $\theta\notin\{\omega,\pi-\omega\}$, then $c-\gamma\notin\sigma(A+B)$.
Proof. Assume the contrary: $c-\gamma\in\sigma(A+B)=\sigma(T(a)+K)$. As $c\neq\gamma$, we have $\lambda_{1}>0$. Since $T(a)+K-\lambda I$ is Fredholm of index zero for $\lambda\in(-\lambda_{1},\lambda_{1})$, the equation
$(T(a)+K-\lambda I)\begin{pmatrix}f_{+}\\ g_{+}\end{pmatrix}=0$
has a nontrivial solution with $f_{+},g_{+}$ in $H^{2}({\mathbf{T}})$ if $\lambda=c-\gamma$. As in the proof of Lemma 5, this equation can be shown to be equivalent to the existence of a constant $\alpha$ such that
$(1+c)[f_{+}]_{0}+(\gamma-c-\lambda)f_{+}(z)+(\delta+sz)g_{+}(z)=0,$
$(\delta z+s)f_{+}(z)+(c-\gamma-\lambda)zg_{+}(z)=\alpha,$
where $f_{+}]_{0}$ denotes te 0th Fourier coefficient of $f_{+}$. For $\lambda=c-\gamma$, this becomes
$(1+c)[f_{+}]_{0}+2(\gamma-c)f_{+}(z)+(\delta+sz)g_{+}(z)=0,$
$(\delta z+s)f_{+}(z)=\alpha.$
If $s\leq\delta$, then $\delta z+s$ has a zero in the closed unit disk and therefore the second equation of the system implies that $\alpha=0$ and that $f_{+}(z)=0$ for all $z$. Inserting this in the first equation, we see that $g_{+}$ must also vanish identically. But this contradicts our assumption, according to which the system must have a nontrivial solution $(f_{+},g_{+})$.
We are left with $s>\delta$. The solution $f_{+}(z)=\alpha/(s+\delta z)$ of the second of these equations is in $H^{2}({\mathbf{T}})$. We have $f_{+}]_{0}=\alpha/s$, and hence the first equation takes the form
$\frac{(1+c)\alpha}{s}+\frac{2(\gamma-c)\alpha}{s+\delta z}+(\delta+sz)g_{+}(z)=0.$
It follows that
$g_{+}(z)=-\frac{\alpha}{\delta+sz}\left[\frac{1+c}{s}+\frac{2(\gamma-c)}{s+\delta z}\right].$
The denominator $\delta+sz$ has the zero $z=-\delta/s$ in the unit disk. Consequently, for $g_{+}$ to be in $H^{2}({\mathbf{T}})$ it is necessary that the term in brackets also vanishes at $z=-\delta/s$. Thus, we must have
$0=\frac{1+c}{s}+\frac{2(\gamma-c)}{s+\delta(-\delta/s)}=\frac{1+c}{s}+\frac{2(\gamma-c)s}{s^{2}-\delta^{2}},$
or equivalently, $(1+c)(s^{2}-\delta^{2})=2s^{2}(c-\gamma)$. Since $s^{2}=1-c^{2}$ and $\delta^{2}=1-\gamma^{2}$, we obtain
$(1+c)(\gamma^{2}-c^{2})=2(1-c^{2})(c-\gamma),$
and as $c\neq-\gamma$, we get $(\gamma+c)=2(1-c)$ and thus $c-\gamma=2$. This contradicts (7) and so proves our claim.    $\square$
Recall that, by Lemma 5,
$),\displaystyle s\leq\delta\;\>\Longleftrightarrow\;\>\theta\in(0,\omega]\cup[\pi-\omega,\pi)\;\>\Longleftrightarrow\;\>c-\gamma\notin\sigma(T({\tilde{a}}$)),
$),\displaystyle s>\delta\;\>\Longleftrightarrow\;\>\theta\in(\omega,\pi-\omega)\;\>\Longleftrightarrow\;\>c-\gamma\in\sigma(T({\tilde{a}}$)),
Theorem 9.
If $A,B$ are given by (10), then $\sigma(A+B)=\Lambda$ for $s\leq\delta$ and $\sigma(A+B)=\Lambda\setminus\{c-\gamma\}$ for $s>\delta$.
Proof. We have $)\Lambda=\sigma(A+B)\cup\sigma(T({\tilde{a}}$)) and $A+B=T(a)+K$. If $s\leq\delta$, then Lemma 5 shows that $\sigma(A+B)=\Lambda$. For $s>\delta$, we have, again by Lemma 5, $\sigma(A+B)=\Lambda$ if $c-\gamma\in\sigma(T(a)+K)$ and $\sigma(A+B)=\Lambda\setminus\{c-\gamma\}$ if $c-\gamma\notin\sigma(T(a)+K)$. But Lemma 8 says that $c-\gamma$ is never in $\sigma(T(a))+K)$ for $s>\delta$, which implies that $\sigma(A+B)=\Lambda\setminus\{c-\gamma\}$ in this case.    $\square$
Theorem 10.
Let $A,B$ be given by (10). If $0<\omega\leq\pi/2$, then
$\varrho([A,B])=\left\{\begin{array}[]{lll}2\sin(\theta+\omega)&\mbox{for}&0<\theta\leq\pi/2-\omega,\\ 2&\mbox{for}&\pi/2-\omega\leq\theta\leq\pi/2+\omega,\\ 2|\sin(\theta-\omega)|&\mbox{for}&\pi/2+\omega\leq\theta<\pi,\end{array}\right$.
and if $\pi/2\leq\omega<\pi$, then
$\varrho([A,B])=\left\{\begin{array}[]{lll}2|\sin(\theta-\omega)|&\mbox{for}&0<\theta\leq\omega-\pi/2,\\ 2&\mbox{for}&\omega-\pi/2\leq\theta\leq 3\pi/2-\omega,\\ 2\sin(\theta+\omega)&\mbox{for}&3\pi/2-\omega\leq\theta<\pi.\end{array}\right$.
Figure 4. The spectra of $P_{200}(A+B)P_{200}$ multiplied by $i$ with $A,B$ as in (10) for $\omega=0.3$ (left), $\omega=2.4$ (right), and $\theta=0.1:0.1:3.1$. The solid lines plot $c-\gamma$, the boundaries of $\lambda_{2},-\lambda_{1}]$ and $\lambda_{1},\lambda_{2}]$, and the values $\pm\sqrt{2}$. The vertical lines have the abscissas $\omega$ and $\pi-\omega$.
Proof. Suppose $0<\omega<\pi/2$. In this case the situation is as in the left picture of Figure 4. The curve described by $\lambda_{2}$ intersects the horizontal line at height $\sqrt{2}$ for the $\theta$ given by $2\sin((\theta+\omega)/2)=\sqrt{2}$, that is, for $\theta=\pi/2-\omega$. It is obvious that for $0<\theta<\pi/2-\omega$ the point $\lambda_{0}=\lambda_{2}$ is a point in $\Lambda$ whose square is closest to $2$, and we know that $\lambda_{0}=\lambda_{2}\in\sigma(A+B)$. Thus, for $0<\theta<\pi/2-\omega$ we get
$\varrho([A,B])=\sqrt{\lambda_{2}^{2}(4\!-\!\lambda_{2}^{2})}=\sqrt{4\sin^{2}\!\frac{\theta\!+\!\omega}{2}\!\left(\!4\!-\!4\sin^{2}\!\frac{\theta\!+\!\omega}{2}\right)}=2\sin(\theta+\omega).$
The curve described by $\lambda_{1}$ is at height $\sqrt{2}$ exactly for $\theta=\pi/2+\omega$, and for $\theta$ in $\pi/2-\omega,\pi/2+\omega]$, the point $\lambda_{0}=\sqrt{2}$ lies in $\sigma(A+B)$ and gives $\varrho([A,B])=\sqrt{2(4-2)}=2$. For $\pi/2+\omega<\theta<\pi-\omega$, we have values $c-\gamma$ whose square is closer to $2$ than $(-\lambda_{1})^{2}$; in the left picture of Figure 4, these values occur for $\theta$ from about $2.0$ to about $2.8$. However, for $\theta\in(\omega,\pi-\omega)$ we have $s>\delta$, and hence, by Theorem 9, these points $c-\gamma$ do not belong to $\sigma(A+B)$. Thus, for $\pi/2+\omega<\theta\leq\pi-\omega$, the points $\lambda_{1}$ and $\lambda_{1}$ are those in $\Lambda$ whose squares are closest to $2$. It follows that $\varrho([A,B])=\sqrt{\lambda_{1}^{2}(4\!-\!\lambda_{1}^{2})}$ for these $\theta$, which, as above, gives $\varrho([A,B])=2|\sin(\theta-\omega)|$. Let finally $\theta$ be in $(\pi-\omega,\pi)$. In this case $c-\gamma>-\lambda_{1}$, implying that the square of $c-\gamma$ is closer to $2$ than the square of $\lambda_{1}$, but from Lemma 8 we know that $c-\gamma$ is never in $\sigma(A+B)$, so that again $\varrho([A,B])=\sqrt{\lambda_{1}^{2}(4\!-\!\lambda_{1}^{2})}$.
The case where $\pi/2\leq\omega<\pi$ can be disposed of similarly.    $\square$